OpenStudy (anonymous):

find the value of c mean values theorm? (x-1)(x-2)(x-3) [0,4]

OpenStudy (nowhereman):

You have to calculate the elevation of the secant between (0, f(0)) and (4, f(4)). Then differentiate your function and see where the derivative assumes this value.

OpenStudy (anonymous):

i am find the derivate but my value of c is wrong ... f(b)-f(a) / b-a this is the formula now find the c

OpenStudy (nowhereman):

yes, choose a=0 and b=4 and calculate it.

OpenStudy (anonymous):

x3 cube - 6x2 square +11x -6=0 now find value of c x3-6x2+11x-6=0

OpenStudy (nowhereman):

what is your problem?

OpenStudy (anonymous):

my prb is what is the value of c just..

OpenStudy (nowhereman):

What do you mean by c? You never said what it is!

OpenStudy (anonymous):

its a formula ok f(c)=f(b)_ f(a) /b-a not find the derivate and put the values

OpenStudy (nowhereman):

You mean \[f'(c) = \frac{f(b) - f(a)}{b-a}\] and what is the problem with it? Just insert all the known variables, do the differentiation and solve the resulting quadratic equation!

OpenStudy (anonymous):

iam put the all values f(b)-f(a)/ b-a. what is the answer using quarditic equation

OpenStudy (nowhereman):

So what quadratic equation have you got?

OpenStudy (anonymous):

i have to values of c ..2-2\[\sqrt{15}\]

OpenStudy (anonymous):

2-2 ander root 15 2+2 under root 15

OpenStudy (nowhereman):

You should check your calculation, the correct solution is \[c = 2 \pm \frac 2 3 \sqrt{3}\]